2021-05-26

From Ciswikidb
Jump to: navigation, search

Meeting Information

Connecting

To connect to the meeting, please do the following: Time: Wednesday at 9:00 AM EST.

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/946721076.
  2. To join via Polycom room system go to the IP Address: 199.48.152.152 (bjn.vc) and enter the meeting ID: 946721076.
  3. To join via phone, use one of the following numbers and the Conference ID: 946721076.
    • Dial Either +1 408 740 7256 or +1 888 240 2560 for US or Canada

Agenda

Schedule

Long Term Schedule: Gantt (Outstanding tasks)

  1. Supplementary Shielding Installation: Mid-June
  2. Beamline Commissioning / Test Plans / Shielding Verification
  3. 7/26/21 Start Hot Check-out
  4. Commission new beam line and execute experiment
  5. Science!!!

New items for discussion:

Short term schedule update:
This week:

  • Area up to transition after solenoid under vacuum. RGA leak check completed.
  • Downstream viewer / faraday cup assembly under vacuum. RGA leak check completed.
  • Last 4 5/8" bpm pulled from elevated line & installed. (One: failed insertion measurements & w/ John Musson, One: stripped tap hole).
  • New valve closed & middle diagnostic cross connected to upstream stuff on 2.75" flanges using bellows & available nipple with 4 5/8" flanges. NEG heated with 1A to drive out H2O. Pump cart is online & pumping @ 1500 Hz (i.e., good). Downstream spool piece + flange also installed. All but ~2 feet of beam line now under vacuum and leak checked.
  • Vacuum pumping down well, but light bake may be required once we have completed installation.
  • Trim cards being pulled by EES-DC for removed magnets on elevated line.
  • Roof tiles off for HDIce & Dilution refrigerator craned out (5/25). Roof tiles replaced again this morning (5/26).
  • Still more vacuum components dropped off to SRF for cleaning. Should probably order some 4 5/8" flanges...

Still to do:

  • 40 L/s ion pump is attached to downstream side of valve on elevated line via a right angle valve & T. When pump cart is available, we need to get this ion pump ON. That puts elevated beam line in safe state, no accidents if valve were to OPEN (it's closed with controls disconnected).
  • Last spool piece [4.5" flange x 24.5" long x 4 5/8" flange] under fabrication. (Note: Pump cart w/ RGA suffering intermittent leak; Matt isolated to 11 L/s pump & 4-way cross joint - 4-way cross new, so swap out 11 L/s pump if it's an issue. One is staged on shelf in cave.) To install:
    1. Ion pumps OFF
    2. Purge with GN2 from both sides, one side through pump cart, the other side through kwik flange with 1/4" compression port.
    3. Use other pump cart from vacuum lab TL1137 (attached to the LIGO tube station) to vent from both sides, with GN2. Wait until Matt's return if desired.
    4. One more pump down and RGA leak check
    5. Add the thin window
    6. Heat / bake beam line
    7. Activate NEGs
    8. Switch Wastewater line vacuum ion pump HV from local supplies to house power / UHV supplies.
    9. Make UED updates as appropriate for vacuum system.
  • Install target mechanism
  • Add Mu metal over bellows on wastewater line.
  • Gun work progressing well:
    1. HV chamber evaluation - plan to bake to evaporate lubricant & pump it off, install photocathode, and bake again.
    2. Sb source laser aperture added in depo chamber; aligned to laser path.
    3. Mask alignment examined.
    4. Gun rebuild schedule
    5. Minor SF6 leak in HV chamber identified; repair later...


Next Week:

  • Build 80/20 support structure; Pb table components awaiting delivery.
  • Wrap up downstream vacuum work
  • EES may start cabling downstream section & making UED updates.
  • Break vacuum in FC3 region
  • Ready to install Al base plates for FC3 region rework as well as solenoid stand. FC3 plate awaiting machining.

2 weeks:

  • Mount bottle rack in Labyrinth; remove gas bottles to rack.
  • Unistrut mounting for cables - Neal will do when available
  • Pull Heliax from Spectrum analyzer/Control room >> Buncher
  • 4th terminal system in Control room under assembly/rebuild
  • Items from To-do list...

Engineering Other business

Reports on Open Action Items:

Action Items:

  • Formal run plan (McCaughan)

Back to Main Meeting Minutes Page